compare the stability of the following compounds,
(i) (a) ph-ch2(free radical)  (b)(ch3)3c(free radical)  (c)ch3-ch(free radical)-ch3  (d)ch2=ch-ch2(free radical)
 
(ii)(a) ch2=ch(free radical)  (b)ph-ch2(free radical)   (c)ph(free radical)  (d)ch(triple bond)c(free radical)  (e)ch3-ch2(free radical)     
Kindly give me DETAILED EXPLANATION.

Grade:12

5 Answers

sufi digital
9 years ago
1) (ch3)3c(free radical) is more stable than ch3-ch(free radical)-ch3 as greater the number of alkyl groups attached to the C-atom with odd electrons more stable is the free radical. this is because of hyperconjugation.  ph-ch2(free radical) is more stable than both these radicals due to resonance. ch2=ch-ch2(free radical) is least stable.
thus the order of stability is ph-ch2(free radical)>(ch3)3c(free radical)>ch3-ch(free radical)-ch3>ch2=ch-ch2(free radical).
​2) ph(free radical) is least stable. ph-ch2(free radical) is most stable.
ch2=ch(free radical)(allyl) is more stable than ch3-ch2(free radical)(primary). 
order of stability is ph-ch2(free radical)>ch2=ch(free radical)>ch3-ch2(free radical)>ph(free radical)
i don’t know about ch(triple bond)c(free radical).
stability of free radicals depend on two factors hyperconjugation and resonance.
hope this helps!!!
vineet chatterjee
9 years ago
but the answer given is ph-ch2(free radical)>ch2=ch-ch2(free radical)>(ch3)3c(free radical)>ch3-ch(free radical)-ch3 
 
vineet chatterjee
9 years ago
and for the second question,answer is 
ph-ch2(free radical)>ch3-ch2(free radical)>ch2=ch(free radical)>ph(free radical)>ch(triple bond)c(free radical)
sufi digital
9 years ago
 ch2=ch-ch2(free radical) is an allyl i mistook it for vinyl. allyls are more reactive than primary secondary and tertiary radicals and vinyls are less reactive. ch2=ch(free radical) is a vinyl. i got confused between the two. very sorry for the confusion. 
 
 
sufi digital
9 years ago
actually i meant allyls are more stable than primary secondary and tertiary radicals and vinyl is less stable. allyls are stabalised by resonance but vinyls are not. by mistake i wrote allyls are more reactive. sorry again